A 2.5 kg otter starts from rest at the top of a
muddy incline 86.2 cm long and slides down
to the bottom in 0.50 s.
What net external force acts on the otter
along the incline?
Answer in units of N.
1
Expert's answer
2019-12-19T13:01:48-0500

First of all, we determine the acceleration "a" the otter descended. To do this, we write out the formula for uniformly accelerated motion "s=a\cdot \frac{t^2}{2}" . From this formula we find the acceleration

"a=2\cdot \frac{s}{t^2}=2\cdot \frac{0.862 m}{0.50^2 s^2}=6.896\ m/s^2" . The net force that caused this acceleration for a body of mass "m=2.5 kg" is determined by Newton’s second law "F=m\cdot a=2.5 kg\cdot 6.896 m/s^2=17.24 N"

Answer: The net external force acts on the otter along the incline is "17.24 N"